Use the elimination method to solve the system of equations.
a+7b=2
a+4b=-1
These are the chooses of answers.
A. There is one solution, it is____.(Simplify your answer. Type an ordered pair)
B. There are infinitely many solutions.
C. There is no solutions.

Found 2 solutions by Alan3354, lynnlo:
Answer by Alan3354(69443)

Use the elimination method to solve the system of equations.
a+7b=2
a+4b=-1
---------- Subtract

Answer by lynnlo(4176)

a+7b=2
a+4b=-1
(A)
a=-5
B=1
THIS PROBLEM IS INDEPENDENT
ELIMINATION IS SUBTRACTING